# Least multivariate Chebyshev polynomials on diagonally determined domains

### v1

@@ Abstract.
We consider a new multivariate generalization of the classical monic (univariate) Chebyshev polynomial that minimizes the uniform norm on the interval $[-1,1]$. Let $\Pi^{*}_{n}$ be the subset of polynomials of degree at most $n$ in $d$ variables, whose homogeneous part of degree $n$ has coefficients summing up to $1$. The problem is determining a polynomial in $\Pi^{*}_{n}$ with the smallest uniform norm on a domain $\Omega$, which we call a least Chebyshev polynomial (associated with $\Omega$). Our main result solves the problem for $\Omega$ belonging to a non-trivial class of domains, defined by a property of its diagonal, and establishes the remarkable result that a least Chebyshev polynomial can be given via the classical, univariate, Chebyshev polynomial. In particular, the solution can be independent of the dimension. The result is valid for fairly general domains that can be non-convex and highly irregular.
@@ 1. Introduction
of degree $n-1$ is the best polynomial of approximation to $x^{n}$ on $[-1,1]$; that is
@@
The main purpose of this paper is to report our findings on the optimization problem (1.3). It turns out, much to our surprise, that the problem (1.3) can be solved analytically for a fairly general family of domains $\Omega$ in ${\mathbb{R}}^{d}$ for all $d\geq 2$. This family of domains will be referred to as diagonally determined.
@@
The paper is organized as follows. We start by defining and describing several examples of diagonally determined domains in Section 2. Our main results are presented in Section 3, where we describe the least polynomial for a diagonally determined domain. We discuss the dual problem of (1.3) in Section 4 and show that it too has a closed-form solution in the case of diagonally determined sets. We also rephrase this result for the dual problem in the framework of extremal signatures.
2. Diagonally determined domains
@@
We call a set $\Omega\subset\mathbb{R}^{d}$ diagonally determined if the following two conditions hold:
@@
Importantly, Definition 2.1 covers domains that could be non-convex, non-compact, and even highly irregular. Figure 1 gives an example of a non-convex diagonally determined set in ${\mathbb{R}}^{2}$.
@@
Figure 1. Example of a diagonally determined, non-convex set. Here, $\mbox{diag}(\Omega)=[1,3]$, and ${\mathbf{v}}=(1,0)$.
As a first observation, a dilation of a diagonally determined set is again diagonally determined.
@@
If $\Omega\subset\mathbb{R}^{d}$ is diagonally determined with ${\rm diag}(\Omega)=[a,b]$ and vector ${\mathbf{v}}$, then, for any $r>0$, the dilation
@@
is also diagonally determined with ${\rm diag}(r\Omega)=[ra,rb]$ and vector ${\mathbf{v}}$.
@@
We will now show that balls in ${\mathbb{R}}^{d}$ (in any norm) are examples of diagonally determined sets.
@@
Let ${\mathbf{v}}={\mathbf{u}}/\|\mathbf{1}\|$, where ${\mathbf{u}}\in{\mathbb{R}}^{d}$ is ‘dual’ to $\mathbf{1}$, in the sense that $\|{\mathbf{u}}\|_{*}=1$ and $\langle{\mathbf{u}},\mathbf{1}\rangle=\|\mathbf{1}\|$. We then have ${\rm diag}(\Omega)=[-b,b]$ with $b=1/\|\mathbf{1}\|$, $\langle{\mathbf{u}}/\|\mathbf{1}\|,\mathbf{1}\rangle=1$, and $|\langle{\mathbf{u}}/\|\mathbf{1}\|,{\mathbf{x}}\rangle|\leq(1/\|\mathbf{1}\|)\|{\mathbf{u}}\|_{*}\|{\mathbf{x}}\|\leq 1/\|\mathbf{1}\|=b$, so that $\langle{\mathbf{u}}/\|\mathbf{1}\|,{\mathbf{x}}\rangle\in[-b,b]=[a,b]$ for all ${\mathbf{x}}\in\Omega$. ∎
@@
Assume $\Omega\subset{\mathbb{R}}^{d}$ is a subset of a ball in ${\mathbb{R}}^{d}$ in any norm, centered at the origin, and that the diagonal of $\Omega$ coincides with the diagonal of the ball. Then $\Omega$ is a diagonally determined set.
@@
Figure 2. Example of a diagonally determined, non-convex set. Here, $\mbox{diag}(\Omega)=\left[-\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right]$, and ${\mathbf{v}}=\left(\frac{1}{2},\frac{1}{2}\right)$.
@@
The intersection of certain balls in ${\mathbb{R}}^{d}$ with the nonnegative orthant in ${\mathbb{R}}^{d}$ are also diagonally determined, as the next result shows.
@@
Further assume $\Omega\subset{\mathbb{R}}^{d}$ has diagonal $\mathrm{diag}(\Omega)=[0,r/\|\mathbf{1}\|]$ for some $r>0$, and, for all ${\mathbf{x}}\in\Omega$, ${\mathbf{x}}\geq\mathbf{0}$ and $\|{\mathbf{x}}\|\leq r$. Then $\Omega$ is a diagonally determined set with vector ${\mathbf{v}}=\mathbf{u}/\|\mathbf{1}\|$.
@@
The assumption on the norm in Proposition 2.6 is met by, for example, all $\ell^{p}$-norms, and the owl-norm in Example 2.5. As a simple corollary of the proposition, the simplex
@@
is diagonally determined, since it is the intersection of the unit $\ell^{1}$-ball with the nonnegative orthant. See Figure 3 for an illustrative example of Proposition 2.6.
@@
3. Least Chebyshev polynomial for a diagonally determined domain
@@
for the special case where $\Omega$ is diagonally determined.
@@
Let $\Omega$ be diagonally determined with vector ${\mathbf{v}}$ and $\mathrm{diag}(\Omega)=[a,b]$. Then
@@
Moreover, as $\Omega$ is diagonally determined, then for all ${\mathbf{x}}\in\Omega$, one has $\langle{\mathbf{v}},{\mathbf{x}}\rangle\in[a,b]$, and therefore
@@
The minimal value in (3.1) does not dependent explicitly on the dimension $d$, but the diagonal $[a,b]$ may. For example, if $\Omega$ is the unit Euclidean ball in ${\mathbb{R}}^{d}$, then $-a=b=1/\sqrt{d}$.
@@
as shown in [12], which depends on the dimension $d$.
@@
We point out that the requirement $\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a_{\alpha}=1$, is susceptive under linear change of variables other than a translation. For instance, the triangle ^$\hat{\triangle}^{2}=\{\,{\mathbf{x}}:x_{1}\geq 0,x_{2}\geq 0,x_{1}+x_{2}\leq 1\,\}$, is a mirror image of $\triangle^{2}$ under $x_{2}\mapsto 1-x_{2}$. However, under the linear change of variables $x_{2}\mapsto 1-x_{2}$, the leading monomial $a_{0}x_{1}^{2}+a_{1}x_{1}x_{2}+a_{2}x_{2}^{2}$ of a polynomial in $\Pi_{2}^{*}$ becomes, $a_{0}x_{1}^{2}-a_{1}x_{2}x_{2}+a_{2}x_{2}^{2}$, which is no longer an element of $\Pi_{2}^{*}$.
@@ 4. The dual framework and signatures
In this section we consider the dual problem of (1.3). Our goal is to show that, in the special case of diagonally determined domains, the dual problem has a closed-form solution. Moreover, we will construct a dual solution that is atomic (discrete), and supported on $n+1$ points.
@@ 4.1. The dual problem
Let $\Omega\subset{\mathbb{R}}^{d}$, and consider the dual problem of (1.3), namely
@@
One may derive this result directly through conic linear programming duality theory, e.g. [11, Proposition 2.9], and we will omit the proof here.
@@
Recall the strong duality result for the classical Chebyshev approximation problem as given in the paper by Rivlin and Shapiro [10, Corollary 2]. [^1] In [10, Corollary 2], the objective is in fact given as $\sup|L(f)|$, but the absolute value may be omitted w.l.o.g., since $L$ is feasible for the dual problem if and only if $-L$ is feasible.
@@ 4.2. Atomic dual solutions and signatures
We first recall a classical interpolation formula for linear functionals, as given in [10, Corollary 3].
@@
The first statement follows immediately from Theorem 4.4, using $V=\Pi_{n}^{d}$. It remains to show that the points $\omega_{1},\ldots,\omega_{r}$ are extremal points of any optimal solution $P^{*}$ to problem (1.3). Now let $L^{*}$ denote an optimal solution of problem (4.1), so that
@@
The extremal points in Corollary 4.5 are usually called the support of an extremal signature, defined along the line of [9, Section 2.2] as follows.
@@
This is essentially a reformulation of the well-known characterization of best approximation by polynomials in terms of the extremal signature ([9, Theorem 2.6]) for the dual problem.
4.3. Signatures for diagonally determined domains
For the special case where $\Omega$ is a diagonally determined set, we may now infer information about an optimal signature from Theorem 3.1 and Proposition 4.7. The idea of the proof is to construct an atomic optimal solution for the dual problem (4.1).
@@ 4.3. Signatures for diagonally determined domains
Let $\Omega$ be a diagonally determined set with $\mathrm{diag}(\Omega)=[a,b]$ and vector ${\mathbf{v}}$. Then, an optimal signature for problem (1.3) is defined by a set $S$, of $n+1$ points
@@
Next, we will construct an atomic solution of the dual problem (4.1), that will lead to the required optimal signature, by Proposition 4.7.
@@
$L^{*}=\frac{1}{n+1}\sum_{i=1}^{n+1}(-1)^{i+1}\lambda_{i}L_{\omega_{i}}.$
@@
(4.7) $L^{*}(T_{k})=\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}T_{k}(\xi_{j})=0\mbox{ if $k<n$.}$
@@
$L^{*}(p)$ $=$ $\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}T_{k_{1}}(\xi_{j})T_{k_{2}}(\xi_{j})$
$=$ $\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}\frac{1}{2}[T_{|k_{1}-k_{2}|}(\xi_{j})+T_{k_{1}+k_{2}}(\xi_{j})]$
$=$ $\frac{1}{2}\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}T_{|k_{1}-k_{2}|}(\xi_{j})+\frac{1}{2}\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}T_{k_{1}+k_{2}}(\xi_{j})=0,$

### v2

# Least multivariate Chebyshev polynomials on diagonally-determined domains

@@ Abstract.
We consider a new multivariate generalization of the classical monic (univariate) Chebyshev polynomial that minimizes the uniform norm on the interval $[-1,1]$. Let $\Pi^{*}_{n}$ be the subset of polynomials of degree at most $n$ in $d$ variables, whose homogeneous part of degree $n$ has coefficients summing up to $1$. The problem is determining a polynomial in $\Pi^{*}_{n}$ with the smallest uniform norm on a domain $\Omega$, which we call a least Chebyshev polynomial (associated with $\Omega$). Our main result solves the problem for $\Omega$ belonging to a non-trivial class of domains that we call diagonally-determined, and establishes the remarkable result that a least Chebyshev polynomial can be given via the classical, univariate, Chebyshev polynomial. In particular, the solution can be independent of the dimension. Diagonally-determined domains include centered balls in $\mathbb{R}^{d}$ in any norm, but can be non-convex and highly irregular. We also introduce a computational procedure, based on semidefinite programming hierarchies, to detect if a given semi-algebraic set is diagonally-determined.
@@ 1. Introduction
of degree $n-1$ ($n\in\mathbb{N}$) is the best polynomial of approximation to $x^{n}$ on $[-1,1]$; that is
@@
The main purpose of this paper is to report our findings on the optimization problem (1.3). It turns out, much to our surprise, that the problem (1.3) can be solved analytically for a fairly general family of domains $\Omega$ in ${\mathbb{R}}^{d}$ for all $d\geq 2$. This family of domains will be referred to as diagonally-determined.
@@
The paper is organized as follows. We start by defining and describing several examples of diagonally-determined domains in Section 2. Our main results are presented in Section 3, where we describe the least polynomial for a diagonally-determined domain. We discuss the dual problem of (1.3) in Section 4 and show that it too has a closed-form solution in the case of diagonally-determined sets. We also rephrase this result for the dual problem in the framework of extremal signatures. Finally, we consider the problem of deciding whether a given semi-algebraic set is diagonally determined in Section 5, and propose an algorithmic approach based on semidefinite programming hierarchies. We conclude the paper with a discussion of related problems and open questions in Section 6.
2. Diagonally-determined domains
@@
We call a set $\Omega\subset\mathbb{R}^{d}$ diagonally-determined if the following two conditions hold:
@@
Importantly, Definition 2.1 covers domains that could be non-convex, non-compact, and even highly irregular. Figure 1 gives an example of a non-convex diagonally-determined set in ${\mathbb{R}}^{2}$.
@@
Figure 1. Example of a diagonally-determined, non-convex set. Here, $\mbox{diag}(\Omega)=[1,3]$, and ${\mathbf{v}}=(1,0)$.
As a first observation, a dilation of a diagonally-determined set is again diagonally-determined.
@@
If $\Omega\subset\mathbb{R}^{d}$ is diagonally-determined with ${\rm diag}(\Omega)=[a,b]$ and vector ${\mathbf{v}}$, then, for any $r>0$, the dilation
@@
is also diagonally-determined with ${\rm diag}(r\Omega)=[ra,rb]$ and vector ${\mathbf{v}}$.
@@
We will now show that balls in ${\mathbb{R}}^{d}$ (in any norm), centered at the origin, are examples of diagonally-determined sets.
@@
$\|{\mathbf{x}}\|\leq\||{\mathbf{x}}|\|\quad\forall{\mathbf{x}}\in\mathbb{R}^{d},$
where $|{\mathbf{x}}|=(|x_{1}|,\ldots,|x_{d}|)$, then the vector ${\mathbf{v}}$ may be assumed to be entrywise-nonnegative, denoted by  ${\mathbf{v}}\geq\mathbf{0}$.
@@
Let ${\mathbf{v}}={\mathbf{u}}/\|\mathbf{1}\|$, where ${\mathbf{u}}\in{\mathbb{R}}^{d}$ is ‘dual’ to $\mathbf{1}$, in the sense that $\|{\mathbf{u}}\|_{*}=1$ and $\langle{\mathbf{u}},\mathbf{1}\rangle=\|\mathbf{1}\|$. We then have ${\rm diag}(\Omega)=[a,b]$ with $a=-1/\|\mathbf{1}\|$ and $b=1/\|\mathbf{1}\|$, $\langle{\mathbf{u}}/\|\mathbf{1}\|,\mathbf{1}\rangle=1$, and $|\langle{\mathbf{u}}/\|\mathbf{1}\|,{\mathbf{x}}\rangle|\leq(1/\|\mathbf{1}\|)\|{\mathbf{u}}\|_{*}\|{\mathbf{x}}\|\leq 1/\|\mathbf{1}\|=b$, so that $\langle{\mathbf{u}}/\|\mathbf{1}\|,{\mathbf{x}}\rangle\in[a,b]$ for all ${\mathbf{x}}\in\Omega$.
[…]
$\|{\mathbf{u}}\|_{*}\geq\||{\mathbf{u}}|\|_{*}\quad\forall{\mathbf{u}}\in\mathbb{R}^{d}.$
Thus, the vector ${\mathbf{u}}$ ‘dual’ to $\mathbf{1}$ in the above construction is component-wise nonnegative. Indeed, if not, then
$1=\|{\mathbf{u}}\|_{*}\geq\||{\mathbf{u}}|\|_{*}\geq\langle|{\mathbf{u}}|,\mathbf{1}/\|\mathbf{1}\|\rangle>\langle{\mathbf{u}},\mathbf{1}/\|\mathbf{1}\|\rangle=1,$
a contradiction. Subsequently, since ${\mathbf{v}}={\mathbf{u}}/\|\mathbf{1}\|$, we have ${\mathbf{v}}\geq\mathbf{0}$, as required. ∎
@@
Assume $\Omega\subset{\mathbb{R}}^{d}$ is a subset of a ball in ${\mathbb{R}}^{d}$ in any norm, centered at the origin, and that the diagonal of $\Omega$ coincides with the diagonal of the ball. Then $\Omega$ is a diagonally-determined set.
@@
Figure 2. Example of a diagonally-determined, non-convex set. Here, $\mbox{diag}(\Omega)=\left[-\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right]$, and ${\mathbf{v}}=\left(\frac{1}{2},\frac{1}{2}\right)$.
@@
The intersection of certain balls in ${\mathbb{R}}^{d}$ with the nonnegative orthant in ${\mathbb{R}}^{d}$ is also diagonally-determined, as the next result shows.
@@
Further assume $\Omega\subset{\mathbb{R}}^{d}$ has diagonal $\mathrm{diag}(\Omega)=[0,r/\|\mathbf{1}\|]$ for some $r>0$, and, for all ${\mathbf{x}}\in\Omega$, ${\mathbf{x}}\geq\mathbf{0}$ and $\|{\mathbf{x}}\|\leq r$. Then $\Omega$ is a diagonally-determined set with vector ${\mathbf{v}}=\mathbf{u}/\|\mathbf{1}\|$.
@@
We note again that the assumption on the norm in Proposition 2.6 is met by, for example, all monotone norms, as shown in the proof of Proposition 2.3.
As a simple corollary of the proposition, the simplex
@@
is diagonally-determined, since it is the intersection of the unit $\ell^{1}$-ball with the nonnegative orthant. See Figure 3 for an illustrative example of Proposition 2.6.
@@
Finally, the diagonally-determined property is not invariant under translation in general. The exception is a translation by a constant multiple of the all-ones vector. We state this observation as a lemma for later use. The proof again follows immediately from Definition 2.1 and is omitted.
Lemma 2.7.
Fix $\alpha\in\mathbb{R}$. Then a given $\Omega\subset\mathbb{R}^{d}$ is diagonally determined with $\mathrm{diag}(\Omega)=[a,b]$ and vector ${\mathbf{v}}$ if and only if its translation $\Omega+\alpha\mathbf{1}$ is diagonally determined with diagonal $[a+\alpha,b+\alpha]$ and vector ${\mathbf{v}}$.
3. Least Chebyshev polynomial for a diagonally-determined domain
@@
for the special case where $\Omega$ is diagonally-determined.
@@
Let $\Omega$ be diagonally-determined with vector ${\mathbf{v}}$ and $\mathrm{diag}(\Omega)=[a,b]$. Then
@@
Moreover, as $\Omega$ is diagonally-determined, then for all ${\mathbf{x}}\in\Omega$, one has $\langle{\mathbf{v}},{\mathbf{x}}\rangle\in[a,b]$, and therefore
@@
The minimal value in (3.1) does not depend explicitly on the dimension $d$, but the diagonal $[a,b]$ may. For example, if $\Omega$ is the unit Euclidean ball in ${\mathbb{R}}^{d}$, then $[a,b]=[-1/\sqrt{d},1/\sqrt{d}]$.
@@
as shown in [18] (see also [15]), which depends on the dimension $d$.
@@
Example 3.4.
Returning to Example 2.5, where $\Omega$ is the unit ball for the owl norm, we have $[a,b]=[-1/W,1/W]$ and ${\mathbf{v}}={\mathbf{w}}/W$, so that, by Theorem 3.1, the least Chebyshev polynomial of degree $n$ becomes
${\mathbf{x}}\mapsto P^{*}({\mathbf{x}})=\left(\frac{1}{W}\right)^{n}\frac{1}{2^{n-1}}T_{n}\left(\langle{\mathbf{w}},{\mathbf{x}}\rangle\right).$
We point out that the requirement $\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a_{\alpha}=1$, is susceptible to an affine change of variables other than a translation. For instance, the triangle ^$\hat{\triangle}^{2}=\{\,{\mathbf{x}}:x_{1}\geq 0,x_{2}\geq 0,x_{1}+x_{2}\leq 1\,\}$, is a mirror image of $\triangle^{2}$ under $x_{2}\mapsto 1-x_{2}$. However, under the affine change of variables $x_{2}\mapsto 1-x_{2}$, the leading monomial $a_{0}x_{1}^{2}+a_{1}x_{1}x_{2}+a_{2}x_{2}^{2}$ of a polynomial in $\Pi_{2}^{*}$ becomes, $a_{0}x_{1}^{2}-a_{1}x_{1}x_{2}+a_{2}x_{2}^{2}$, which is no longer an element of $\Pi_{2}^{*}$.
Moreover, the extremal problem (1.3) in Definition 1.1 is not rotationally invariant. To see this consider the following three sets in $\mathbb{R}^{2}$, that are rotations of the same line segment:
$\Omega_{1}$ ::=:= $\{(x,x)\;|\;x\in[-1,1]\}$
$\Omega_{2}$ ::=:= $\{(0,x)\;|\;x\in[-\sqrt{2},\sqrt{2}]\}$
$\Omega_{3}$ ::=:= $\{(x,-x)\;|\;x\in[-1,1]\}.$
Of these three sets, only $\Omega_{1}$ is diagonally-determined. Even though the three sets are rotations of each other, the solution to the extremal problem (1.3) is starkly different in each case:
  • For $\Omega_{1}$, Theorem 3.1 provides the solution $(x_{1},x_{2})\mapsto\frac{1}{2^{n-1}}T_{n}(x_{1})$ as a least Chebyshev polynomial, and thus $\inf_{P\in\Pi^{*}_{n}}\,\|P\|_{\Omega_{1}}=\frac{1}{2^{n-1}}$;
  • For $\Omega_{2}$, a least Chebyshev polynomial is clearly given by $(x_{1},x_{2})\mapsto x_{1}^{n}$, so that $\inf_{P\in\Pi^{*}_{n}}\,\|P\|_{\Omega_{2}}=0$;
  • For $\Omega_{3}$, a least Chebyshev polynomial is $(x_{1},x_{2})\mapsto\frac{1}{2^{n}}(x_{1}+x_{2})^{n},$ so that $\inf_{P\in\Pi^{*}_{n}}\,\|P\|_{\Omega_{3}}=0$.
The example shows that the diagonal of $\Omega$ plays a key role in determining the least Chebyshev polynomial, and that it is not only an artefact of the analysis.
@@ 4. The dual framework and signatures
In this section we consider the dual problem of (1.3). Our goal is to show that, in the special case of diagonally-determined domains, the dual problem has a closed-form solution. Moreover, we will construct a dual solution that is atomic (discrete), and supported on $n+1$ points.
@@ 4.1. The dual problem
Let $\Omega\subset{\mathbb{R}}^{d}$ be compact, and consider the dual problem of (1.3), namely
@@
One may derive this result directly through conic linear programming duality theory, e.g. [14, Proposition 2.9], and we will omit the proof here.
@@
Recall the strong duality result for the classical Chebyshev approximation problem as given in the paper by Rivlin and Shapiro [13, Corollary 2]. [^1] In [13, Corollary 2], the objective is in fact given as $\sup|L(f)|$, but the absolute value may be omitted w.l.o.g., since $L$ is feasible for the dual problem if and only if $-L$ is feasible.
@@ 4.2. Atomic dual solutions and signatures
We first recall a classical interpolation formula for linear functionals, as given in [13, Corollary 3].
@@
The first statement follows immediately from Theorem 4.4, applied to an optimal solution of problem (4.1), and using $V=\Pi_{n}^{d}$. It remains to show that the points $\omega_{1},\ldots,\omega_{r}$ are extremal points of any optimal solution $P^{*}$ to problem (1.3). Now let $L^{*}$ denote an optimal solution of problem (4.1), so that
@@
The extremal points in Corollary 4.5 are usually called the support of an extremal signature, defined along the line of [12, Section 2.2] as follows.
@@
This is essentially a reformulation of the well-known characterization of best approximation by polynomials in terms of the extremal signature ([12, Theorem 2.6]) for the dual problem.
4.3. Signatures for diagonally-determined domains
For the special case where $\Omega$ is a diagonally-determined set, we may now infer information about an optimal signature from Theorem 3.1 and Proposition 4.7. The idea of the proof is to construct an atomic optimal solution for the dual problem (4.1).
@@ 4.3. Signatures for diagonally-determined domains
Let $\Omega$ be a diagonally-determined set with $\mathrm{diag}(\Omega)=[a,b]$ and vector ${\mathbf{v}}$. Then, an optimal signature for problem (1.3) is defined by a set $S$, of $n+1$ points
@@
Next, we will construct an atomic solution of the dual problem (4.1), which will lead to the required optimal signature, by Proposition 4.7.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{*}=\frac{1}{n}\sum_{i=1}^{n+1}(-1)^{i+1}\lambda_{i}L_{\omega_{i}}.}$
@@
(4.7) $L^{*}(T_{k})=\frac{1}{n}\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}T_{k}(\xi_{j})=0\quad\mbox{ if $k<n$.}$
@@
$L^{*}(p)$ $=$ $\frac{1}{n}\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}T_{k_{1}}(\xi_{j})T_{k_{2}}(\xi_{j})$
$=$ $\frac{1}{n}\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}\frac{1}{2}[T_{|k_{1}-k_{2}|}(\xi_{j})+T_{k_{1}+k_{2}}(\xi_{j})]$
$=$ $\frac{1}{2n}\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}T_{|k_{1}-k_{2}|}(\xi_{j})+\frac{1}{2n}\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}T_{k_{1}+k_{2}}(\xi_{j})=0,$